Dr. Blumen invested $5000. Part of it was invested in bonds at a rate of 6% and the rest was invested in a money market at a rate of 7.5%. If the annual interest is $337.50, how much did she invest in bonds?
Answer by mananth(16949)

Bonds 6.00% per annum ---x
Money market 7.50% per annum ---y

x+y= 5000 ------------------------1
6.00% x+7.50% y=337.5
Multiply by 100
6x+7.5y= = 33750.00 --------2
Multiply (1) by -6
we get
-6x-6y= -30000.00
Add this to (2)
0x+1.5y=3750
divide by 1.5
y = 2500 Bonds 7.50%
Balance 2500 Money market 6.00%
CHECK
2500 --------- 7.50% ------- 187.50
2500 ------- 6.00% ------- 150.00
Total -------------- 337.50